# supplement

### v1

@@ 6 Simulation Study
[…]
@@
Note that HEAT estimation is essentially the first-round output of IteHEAT. Therefore, we present the performance of the IteHEAT estimators $\widetilde{\Omega}^{(\cdot,t)}$ for $t=0,1,2,\ldots$, where $\widetilde{\Omega}^{(\cdot,1)}$ represents the HEAT estimator, and $\widetilde{\Omega}^{(\cdot,0)}:=\widehat{\Omega}^{(\cdot)}$ represents the individual local estimator in (4)-(5). The practical determination of regularization parameters $\lambda_{j}^{(m)},j\in[p],m\in[M]$ in Lasso procedures (3) and (22), as well as the shrinkage levels $\lambda_{1,j,k},\lambda_{2,j,k}$ and $\lambda_{1,j,k}^{(t)},\lambda_{2,j,k}^{(t)},t=2,3,\ldots$ in HEAT (12) and IteHEAT (25), is discussed in Sections LABEL:sec:reg_para-LABEL:sec:shrink_para of the supplement. Section LABEL:sec:splitting_effect of the supplement evaluates the effects of data splitting on the proposed IteHEAT and shows that IteHEAT without splitting exhibits superior performance. Therefore, throughout this section, IteHEAT is implemented without data splitting. We employ the soft and SCAD thresholding functions, and denote the corresponding methods by HEAT & IteHEAT_Soft and HEAT & IteHEAT_SCAD, respectively. For illustration purpose, we set $n_{0}=400,p\in\{100,200\}$ and $M\in\{5,20\}$.
[…]
@@ 7 Real Data Analysis
In this section, we explore the version 8 bulk tissue RNA-seq count data in Genotype-Tissue Expression (GTEx) project. This dataset collects samples from 54 tissues across nearly 1000 donors, and can be accessed publicly at https://www.gtexportal.org. We focus on nine tissues with the most frequently collected samples: Adipose-Subcutaneous, Lung, Muscle-Skeletal, Skin-Sun Exposed, Skin-Not Sun Exposed, Thyroid, Nerve-Tibial, Artery-Tibial and Whole Blood (GTEx-Consortium,, 2020), and then employ our proposed methodology to conduct differential analysis for expression networks. Gender differences in immune diseases are widely acknowledged (Moroni et al.,, 2012; Ngo et al.,, 2014). Thus, we investigate gender heterogeneity within each tissue separately, focusing on the gene set related to antigen processing and presentation (GO:0019882), which contains 117 genes. Consequently, we have $M=2$ across tissues. The sample sizes in male datasets range from 395 to 543, while those in female datasets range from 183 to 260.
@@
For each tissue analysis, we employ edgeR (Robinson et al.,, 2010) to preprocess the RNA-seq count datasets. This includes low-expressed gene filtering, normalization, and log-counts-per-million transformation. After preprocessing, the numbers of remaining genes range from 100 to 111. Due to the limited number of integrated datasets ($M=2$), we implement HEAT with SCAD thresholding separately for each tissue. We then extract estimated network edges that exhibit gender heterogeneity. The top ten genes showing the highest connectivity within heterogeneous edges are illustrated in Figure 3. The results reveal that the gene KDM5D exhibits the highest degree of heterogeneity in connectivity. Additionally, heterogeneous associations related to the HLA gene family are observed in several tissues. Notably, KDM5D, located on the Y chromosome, is recognized as a tumor suppressor (Ohguchi and Ohguchi,, 2022), and HLA genes are associated with various immune diseases (Ngo et al.,, 2014). These findings of gender-related genetic association heterogeneity support the validity of the proposed methodology.
8 Conclusion and Discussion
In this article, we focus on distributed integrative estimation for high-dimensional precision matrices under heterogeneity and introduce the HEAT and IteHEAT methods. These proposals achieve both communication and computation efficiency while maintaining the statistical optimality. Importantly, the ideas underlying HEAT and IteHEAT are not limited to precision matrix estimation alone, and can be extended to various models with unbiased estimators that can be facilitated by modern inference theories (e.g., Chernozhukov et al.,, 2018).
In conclusion, we raise two questions for future research. First, while simulations demonstrate the superiority of IteHEAT without data splitting, a rigorous theoretical guarantee is beneficial to validate its performance. Second, the current heterogeneity modelling in HEAT and IteHEAT may be sensitive to highly adversarial or contaminated datasets. Therefore, developing robust methodologies for distributed integration warrants further exploration.

### v2

@@ 6 Simulation Study
In this section, we evaluate the simulation performance of the proposed HEAT and IteHEAT approaches, and compare them with several joint estimation procedures that have access to the full distributed datasets as oracle baselines. The data generation procedure is detailed in Section LABEL:sec:addition_data_generation of the supplement and is briefly described as follows. To construct $\Omega^{(\cdot)}$, two graph structures are considered: the Erdös-Rényi random graph and the banded graph, and the heterogeneity pattern is controlled by a parameter $\texttt{hete\_ratio}\in[0,1]$, which approximately represents the ratio $\frac{s_{2}}{s_{1}+s_{2}}$. Then we generate $X_{i}^{(m)}\sim\text{N}(0,(\Omega^{(m)})^{-1})$ for $i\in[n_{m}]$ and $m\in[M]$, where the sample sizes are independently determined by $n_{m}=n_{0}+\lceil\frac{n_{0}}{25}\text{N}(0,1)\rceil$, with $n_{0}$ being a preassigned value. To evaluate the estimation errors, we consider the loss $L_{1,r}$ in (17) with matrix 1-norm $\|L_{1,r}\|_{1}$ and Frobenius norm $\|L_{1,r}\|_{\text{F}}^{2}/p$, denoted by L_1 and Frobenius respectively. We present the results with $r=1$. All simulation results are based on 256 independent replications unless stated otherwise.
@@
For our numerical experiments, we implement HEAT and IteHEAT using both soft and SCAD thresholding functions for illustration. Note that HEAT estimation is essentially the first-round output of IteHEAT. Therefore, we present the performance of the IteHEAT estimators $\widetilde{\Omega}^{(\cdot,t)}$ for $t=0,1,2,\ldots,10$, where $\widetilde{\Omega}^{(\cdot,1)}$ represents the HEAT estimator, and $\widetilde{\Omega}^{(\cdot,0)}:=\widehat{\Omega}^{(\cdot)}$ represents the individual local estimator in (4)-(5). We refer to the corresponding methods as HEAT & IteHEAT_Soft and HEAT & IteHEAT_SCAD. In addition, we evaluate IteHEAT that employs a data-driven stopping rule for iterations, which is described in Section LABEL:sec:addition_num_iteration of the supplement, and we denote the methods by IteHEAT_Soft (stop) and IteHEAT_SCAD (stop). The practical determination of regularization parameters $\lambda_{j}^{(m)},j\in[p],m\in[M]$ in Lasso procedures (3) and (23), as well as the shrinkage levels $\lambda_{1,j,k},\lambda_{2,j,k}$ and $\lambda_{1,j,k}^{(t)},\lambda_{2,j,k}^{(t)},t=2,3,\ldots$ in HEAT (12) and IteHEAT (27), is discussed in Sections LABEL:sec:addition_reg_para-LABEL:sec:addition_shrink_para of the supplement. For both HEAT and IteHEAT, we employ the entry-dependent adaptive shrinkages, and numerical comparisons between universal and adaptive shrinkages are provided in Section LABEL:sec:addition_compare_shrinkage of the supplement. For IteHEAT, we set the number of splits to $H=10$; see Section LABEL:sec:addition_split of the supplement for further discussion. In the following, we set $n_{0}=400,p\in\{100,200\}$ and $M\in\{5,20\}$.
Figure 1: Comparisons of HEAT & IteHEAT, IteHEAT (stop) and NodewiseLasso_Full under homogeneous settings where $\texttt{hete\_ratio}=0$.
We first examine the homogeneous settings where $\texttt{hete\_ratio}=0$, indicating that $s_{2}=0$ and all datasets share the same distribution. We compare HEAT & IteHEAT and IteHEAT (stop) with a competing oracle method NodewiseLasso_Full, which estimates the homogeneous precision matrix using the node-wise Lasso method introduced in Section 3.1 based on the full dataset $X_{i}^{(m)},i\in[n_{m}],m\in[M]$. The results are displayed in Figure 1. Overall, HEAT & IteHEAT_Soft is comparable to NodewiseLasso_Full, while HEAT & IteHEAT_SCAD demonstrates significant advantages over HEAT & IteHEAT_Soft and NodewiseLasso_Full. Furthermore, the performance of IteHEAT (stop) closely matches the best iteration of HEAT & IteHEAT, which validates the effectiveness of our proposed stopping criterion. These results indicate that the proposed HEAT and IteHEAT exhibit competitive performance in homogeneous settings, demonstrating the validity and statistical accuracy in distributed integration.
Figure 2: Comparisons of HEAT & IteHEAT, IteHEAT (stop), JGL_Full and HGSL_Full under completely heterogeneous settings where $\texttt{hete\_ratio}=1$.
Next, we compare the proposed methods with two alternative approaches under heterogeneity: Joint Graphical Lasso in Danaher et al., (2014) and Heterogeneous Group Square-Root Lasso in Ren et al., (2019), denoted by JGL_Full and HGSL_Full, respectively. Both methods utilize the full distributed datasets and employ the group Lasso penalty for estimation under shared supports. Thus, to ensure a fair comparison, we focus on the completely heterogeneous settings where $\texttt{hete\_ratio}=1$. For JGL_Full, the regularization involves two parameters: $\lambda_{1}$ and $\lambda_{2}$. The selection of $\lambda_{2}$ follows the same criterion (LABEL:eq:joint_criterion) as HEAT and IteHEAT described in Section LABEL:sec:addition_shrink_para of the supplement, while $\lambda_{1}$ is set to 0 to suit the current settings; see Danaher et al., (2014) for more details. For HGSL_Full, the regularization parameter is determined by the tuning-free simulation strategy suggested in Ren et al., (2019). However, due to high computational costs, the performance of HGSL_Full is evaluated based on only 64 replications for $(M,p)=(20,100)$, and it is not implemented for $(M,p)=(20,200)$. The results are displayed in Figure 2. It is clear that JGL_Full exhibits inferior performance compared to all the other methods. While there is a noticeable gap between HEAT & IteHEAT_Soft and HGSL_Full, the proposed HEAT & IteHEAT_SCAD significantly outperforms HGSL_Full, even with $t=1$ in most settings. Besides, consistent with the results in homogeneous settings, IteHEAT (stop) achieves performance comparable to the fully iterated HEAT & IteHEAT. These results further highlight the advantage of HEAT and IteHEAT in terms of statistical accuracy and flexibility in modelling heterogeneity.
@@ 7 Real Data Analysis
In this section, we explore the version 8 bulk tissue RNA-seq count data in Genotype-Tissue Expression (GTEx) project. This dataset collects samples from 54 tissues across nearly 1000 donors, and can be accessed publicly at https://www.gtexportal.org. We focus on nine tissues with the most frequently collected samples (GTEx-Consortium,, 2020): Adipose-Subcutaneous, Lung, Muscle-Skeletal, Skin-Sun Exposed, Skin-Not Sun Exposed, Thyroid, Nerve-Tibial, Artery-Tibial and Whole Blood, and then employ our proposed methodology to conduct differential analysis for expression networks. Gender differences in immune diseases are widely acknowledged (Moroni et al.,, 2012; Ngo et al.,, 2014). Thus, we investigate gender heterogeneity within each tissue separately, focusing on the gene set related to antigen processing and presentation (GO:0019882), which contains 117 genes. Consequently, we have $M=2$ across tissues. The sample sizes in male datasets range from 395 to 543, while those in female datasets range from 183 to 260.
@@
For each tissue analysis, we employ edgeR (Robinson et al.,, 2010) to preprocess the RNA-seq count datasets. This includes low-expressed gene filtering, normalization, and log-counts-per-million transformation. After preprocessing, the numbers of remaining genes range from 100 to 111. Due to the limited number of integrated datasets ($M=2$), we implement HEAT with SCAD thresholding and entry-dependent shrinkage levels separately for each tissue. We then extract estimated network edges that exhibit gender heterogeneity. The top ten genes showing the highest connectivity within heterogeneous edges are illustrated in Figure 3. The results reveal that the gene KDM5D exhibits the highest degree of heterogeneity in connectivity. Additionally, heterogeneous associations related to the HLA gene family are observed in several tissues. Notably, KDM5D, located on the Y chromosome, is recognized as a tumor suppressor (Ohguchi and Ohguchi,, 2022), and HLA genes are associated with various immune diseases (Ngo et al.,, 2014). These findings of gender-related genetic association heterogeneity support the validity of the proposed methodology.
8 Conclusions and Discussions
In this article, we focus on distributed integrative estimation for high-dimensional precision matrices under heterogeneity and introduce the HEAT and IteHEAT methods. These proposals achieve both communication and computation efficiency while maintaining the statistical optimality. More importantly, the ideas underlying HEAT and IteHEAT are not limited to precision matrix estimation alone, and can be extended to various models with unbiased local estimators that can be facilitated by modern inference methods and theories (e.g., van de Geer et al.,, 2014; Chernozhukov et al.,, 2018, 2022).
In conclusion, we raise several problems for future research. First, the IteHEAT in Algorithm 5, which enhances the accuracy and optimality of HEAT in Algorithm 3 through iterative higher-order refinement, requires a restriction on the number of local sample sizes to guarantee the convergence contraction in multi-round communications. Although this requirement is also present in other distributed methods under homogeneity (Wang et al.,, 2017; Jordan et al.,, 2019), developing algorithms free of such restriction is a valuable problem for distributed data integration.
Second, while the proposed HEAT and IteHEAT enjoy communication efficiency as discussed in Remark 2 and Section LABEL:sec:more_discuss_commu of the supplement, adapting them for settings with exact communication constraints remains an open problem. To our knowledge, existing works on communication-constrained sparse estimation have been limited to homogeneous Gaussian mean models (Garg et al.,, 2014; Braverman et al.,, 2016; Han et al.,, 2021; Acharya et al.,, 2023). Extension from homogeneous sparse mean models to heterogeneous sparse matrix models can be non-trivial and is left as future research.
Third, HEAT and IteHEAT ensure the privacy protection of individual-level data by sharing only summary statistics, which aligns with Kohane et al., (2021); Duan et al., (2022); Zhou et al., (2022). Given the recent advances in areas like differential privacy, incorporating quantitative privacy-preserving mechanisms into the developed distributed learning algorithms would also be of interest.
Last, the current heterogeneity modelling in HEAT and IteHEAT may be sensitive to highly adversarial or contaminated datasets. Therefore, developing robust methodologies for distributed integration warrants further exploration.